# Exponential Runge–Kutta methods of collocation type for parabolic equations with time-dependent delay

### v1

@@ 5 Superconvergence
Let the initial value problem (1.1) satisfy Assumptions 2.1-2.2 and 5.1, and consider for its numerical solution the modified ERKC-I method (5.50) whose underlying quadrature rule is of order $s+2$. Let the interpolation approach (5.51) satisfy $\ell_{k}+r_{k}=s+2$ for $k=0,1,\ldots,N-1$ and let the mesh $I_{h}$ be quasi-uniform. If $u:\mathcal{I}^{\mu}\to V$ and $f:\mathcal{I}^{\mu}\to V$ (4.16) are both $s+2$ times differentiable for $\mu=1,2,\ldots,m+1$ with $u^{(s+2)}\in L^{\infty}(0,T;V)$, $A^{\beta}f^{(s)}\in L^{\infty}(0,T;V)$, $f^{(s+1)}\in L^{\infty}(0,T;V)$ and $f^{(s+2)}\in L^{1}(0,T;V)$, then for $h=\max_{1\leq j\leq N}h_{j}$ sufficiently small the error of $U_{n}$ satisfies
@@
Remark 5.3.
The assumption of a quasi-uniform mesh is imposed to ensure that the constant $C$ in (5.52) is mesh-independent. It can instead be replaced by a weaker condition, such as that the step size sequence satisfies $\kappa_{1}\leq h_{j+1}/h_{j}\leq\kappa_{2}$ for all $j$.
@@
$e(t_{n})$ $=\mathrm{e}^{-(t_{n}-t_{0})A}e(t_{0})+\sum_{j=1}^{n}\mathrm{e}^{-(t_{n}-t_{j})A}\bigg{(}\sum_{i=1}^{s}h_{j}b_{i}(-h_{j}A)J_{j-1}e(t_{j-1,i})$ (5.56)
$\quad+h_{j}\sum_{i=1}^{s}b_{i}(-h_{j}A)\Big{(}g(t_{j-1,i},U_{j-1,i},(\Pi^{*}_{h}U)^{\tau}_{j-1,i})-f(t_{j-1,i})-J_{j-1}e(t_{j-1,i})\Big{)}-\delta_{j}(1)\bigg{)}.$
@@
$\delta_{n+1}(c_{i})=h_{n+1}^{s+1}\psi_{i,s+1}(-h_{n+1}A)f^{(s)}(t_{n}^{+})+\tilde{\delta}_{n+1}(c_{i}),$
@@
We can derive the desired bound for the convolution of $\mathrm{e}^{-(t_{n}-t_{j})A}$ with (5.59) in the same manner as presented in [11]. Using (5.57), (5.58), (5.60) and (5.54), we get by applying the discrete Gronwall lemma (Lemma 4.1) the following result: for any mesh point $t_{n}$ belonging to $[0,\xi_{1}]$, it holds
@@
$\qquad\leq\big{\|}g\big{(}t_{j-1,i},U_{j-1,i},(\Pi^{*}_{h}U)^{\tau}_{j-1,i}\big{)}-g\big{(}t_{j-1,i},u(t_{j-1,i}),(\Pi^{*}_{h}U)^{\tau}_{j-1,i}\big{)}-J_{j-1}e(t_{j-1,i})\big{\|}_{X}$
$\qquad\quad+\big{\|}g\big{(}t_{j-1,i},u(t_{j-1,i}),(\Pi^{*}_{h}U)^{\tau}_{j-1,i}\big{)}-f(t_{j-1,i})\big{\|}_{X}$
$\qquad\leq\widetilde{L}\|e(t_{j-1,i})\|_{V}^{2}+L\|(\Pi_{h}^{*}U)^{\tau}_{j-1,i}-u^{\tau}(t_{j-1,i})\|_{V}.$ (5.61)
We solve the error recursion (5.55) for $t_{n}\in(\xi_{\mu},\xi_{\mu+1})$ and consider its estimate in $V$. By using (5.57), (5.58), (5.60) and (5.61), we get by applying the discrete Gronwall lemma (Lemma 4.1) the following result: for any mesh point $t_{n}$ belonging to $[\xi_{\mu},\xi_{\mu+1}]$, it holds
@@ 6 Numerical experiments
We apply standard finite differences with $n=1000$ grid points to discretize the problem (6.62). The resulting products of matrix functions with vectors are computed by the fast Fourier transformation. The convergence rates of ERKC-I and ERKC-C methods at $T=3$ in the $L^{\infty}(\Omega)$ and the $L^{2}(\Omega)$ norms, respectively, are presented in Figures 1 and 2. The numerical results clearly match our theoretical analysis.
@@
We apply standard finite differences with $n=200$ grid points to discretize the problem in each spatial direction. In this example the exact solution is unknown. The reference solution is computed by the ERKC-C method with Gauss collocation points ($s=3$) using the constant step size $h=2^{-11}$. The errors of the ERKC-I and ERKC-C methods in the $L^{\infty}(\Omega)$ and the $L^{2}(\Omega)$ norm at final time $T=3$ are presented in Figures 3 and 4. The figures confirm the theoretical analysis. Moreover, a comparison of computational cost of the ERKC methods with Gauss collocation points ($s=3$) is presented in Table 1. For step sizes $h=2^{-k}$ with $k=3,4,\ldots,9$, the ERKC-I method consistently reduces the CPU time by approximately 12% compared to the ERKC-C method, demonstrating its computational efficiency across a range of step sizes.
@@
for $x\in[0,1]$ and $t\in[0,3]$, subject to homogenous Dirichlet boundary conditions. The initial condition is given by $\phi(t,x)=\mathrm{e}^{t}x(1-x)$ for $t\in[-\frac{1}{2},0]$. The above initial-boundary value problem can be written as an abstract initial value problem with linear operator $A$ defined by $(Au)(x)=-\partial_{xx}u(x)$, and nonlinearity $g$ defined by $g(t,v,w)=v(1-v)-w(1-w)$. The Banach space $X$ can be chosen as $X=L^{2}(\Omega)$ and then $D(A)=H_{0}^{1}(\Omega)\cap H^{2}(\Omega)$. The nonlinearity $g:[0,3]\times V\times V\to X$ is locally Lipschitz-continuous in a strip along the exact solution for $V=H^{\frac{1}{4}}(\Omega)$.
@@
Figure 6: The convergence rates of ERKC-I methods (in the left panel) and modified ERKC-I methods (in the right panel) for (6.65). The errors are measured at $T=3$ in the $L^{\infty}(\Omega)$ norm.

### v3

@@ 5 Superconvergence
Let the initial value problem (1.1) satisfy Assumptions 2.1-2.2 and 5.1, and consider for its numerical solution the modified ERKC-I method (5.13) whose underlying quadrature rule is of order $s+2$. Let the interpolation approach (5.14) satisfy $\ell_{k}+r_{k}=s+2$ for $k=0,1,\ldots,N-1$ and let the step size sequence $\{h_{j}\}_{j=1}^{N}$ satisfy the condition (5.16). If $u:\mathcal{I}^{\mu}\to V$ and $f:\mathcal{I}^{\mu}\to V$ (4.2) are both $s+2$ times differentiable for $\mu=1,2,\ldots,m+1$ with $u^{(s+2)}\in L^{\infty}(0,T;V)$, $A^{\beta}f^{(s)}\in L^{\infty}(0,T;V)$, $f^{(s+1)}\in L^{\infty}(0,T;V)$ and $f^{(s+2)}\in L^{1}(0,T;V)$, then for $h=\max_{1\leq j\leq N}h_{j}$ sufficiently small the error of $U_{n}$ satisfies
@@
$e(t_{n})$ $=\mathrm{e}^{-(t_{n}-t_{0})A}e(t_{0})+\sum_{j=1}^{n}\mathrm{e}^{-(t_{n}-t_{j})A}\bigg(\sum_{i=1}^{s}h_{j}b_{i}(-h_{j}A)J_{j-1}e(t_{j-1,i})$
$\quad+h_{j}\sum_{i=1}^{s}b_{i}(-h_{j}A)\Big(g(t_{j-1,i},U_{j-1,i},(\Pi^{*}_{h}U)^{\tau}_{j-1,i})-f(t_{j-1,i})-J_{j-1}e(t_{j-1,i})\Big)-\delta_{j}(1)\bigg).$ (5.20)
@@
$\delta_{n+1}(c_{i})=h_{n+1}^{s+1}\psi_{i,s+1}(-h_{n+1}A)f^{(s)}(t_{n}^{+})+\widetilde{\delta}_{n+1}(c_{i}),$
@@
We can derive the desired bound for the convolution of $\mathrm{e}^{-(t_{n}-t_{j})A}$ with (5.23) in the same manner as presented in [13]. Using (5.21), (5.22) and (5.18), we get by applying the discrete Gronwall lemma (Lemma 4.1) the following result: for any mesh point $t_{n}$ belonging to $0,\xi_{1}]$, it holds
@@
$\quad\leq\Big\|g\big(t_{j-1,i},U_{j-1,i},u^{\tau}(t_{j-1,i})\big)-f(t_{j-1,i})-\frac{\partial g}{\partial u}\big(t_{j-1,i},u(t_{j-1,i}),u^{\tau}(t_{j-1,i})\big)e(t_{j-1,i})\Big\|_{X}$
$\qquad+\big\|g\big(t_{j-1,i},U_{j-1,i},(\Pi^{*}_{h}U)^{\tau}_{j-1,i}\big)-g\big(t_{j-1,i},U_{j-1,i},u^{\tau}(t_{j-1,i})\big)\big\|_{X}$
$\qquad+\Big\|\frac{\partial g}{\partial u}\big(t_{j-1,i},u(t_{j-1,i}),u^{\tau}(t_{j-1,i})\big)e(t_{j-1,i})-J_{j-1}e(t_{j-1,i})\Big\|_{X}$
$\quad\leq\widetilde{L}\|e(t_{j-1,i})\|_{V}^{2}+L\|(\Pi_{h}^{*}U)^{\tau}_{j-1,i}-u^{\tau}(t_{j-1,i})\|_{V}+Ch\|e(t_{j-1,i})\|_{V}.$ (5.25)
We solve the error recursion (5.19) for $t_{n}\in(\xi_{\mu},\xi_{\mu+1})$ and consider its estimate in $V$. By using (5.21), (5.22), (5.24) and (5.25), we get by applying the discrete Gronwall lemma (Lemma 4.1) the following result: for any mesh point $t_{n}$ belonging to $\xi_{\mu},\xi_{\mu+1}]$, it holds
@@ 6 Numerical experiments
We apply standard finite differences with $n=1000$ grid points to spatially discretize the problem (6.1). The resulting products of matrix functions with vectors are computed by the fast Fourier transformation. The convergence rates of ERKC-I and ERKC-C methods at $T=3$ in the $L^{\infty}(\Omega)$ and the $L^{2}(\Omega)$ norms, respectively, are presented in Figures 1 and 2. The numerical results clearly match our theoretical analysis.
@@
Figure 1: The convergence rates of ERKC-I methods (in the left panel) and ERKC-C methods (in the right panel) for (6.1). The errors are measured at $T=3$ in the $L^{\infty}(\Omega)$ norm. In this example, the two considered approaches produced nearly identical errors. Therefore, the two panels are indistinguishable.
Figure 2: The convergence rates of ERKC-I methods (in the left panel) and ERKC-C methods (in the right panel) for (6.1). The errors are measured at $T=3$ in the $L^{2}(\Omega)$ norm. In this example, the two considered approaches produced nearly identical errors. Therefore, the two panels are indistinguishable.
@@
We apply standard finite differences with $n=200$ grid points to discretize the problem in each spatial direction. In this example the exact solution is unknown. The reference solution is computed by the ERKC-C method with Gauss collocation points ($s=3$) using the constant step size $h=2^{-11}$. The resulting mesh includes the primary discontinuities. The errors of the ERKC-I and ERKC-C methods in the $L^{\infty}(\Omega)$ and the $L^{2}(\Omega)$ norm at final time $T=3$ are presented in Figures 3 and 4. The figures confirm the theoretical analysis. Moreover, a comparison of computational cost of the ERKC methods with Gauss collocation points ($s=3$) is presented in Table 1. For step sizes $h=2^{-k}$ with $k=3,4,\ldots,9$, the ERKC-I method consistently reduces the CPU time by approximately 12% compared to the ERKC-C method, demonstrating its computational efficiency across a range of step sizes.
@@
for $x\in[0,1]$ and $t\in[0,3]$, subject to homogenous Dirichlet boundary conditions. The initial condition is given by $\phi(t,x)=\mathrm{e}^{t}x(1-x)$ for $t\in[-\frac{1}{2},0]$. The above initial-boundary value problem can be written as an abstract initial value problem with linear operator $A$ defined by $(Au)(x)=-\partial_{xx}u(x)$, and nonlinearity $g$ defined by g(t,v,w)=v(1v)w(1w)g(t,v,w)=v(1-v)-w(1-w). The Banach space $X$ can be chosen as $X=L^{2}(\Omega)$ and then $D(A)=H_{0}^{1}(\Omega)\cap H^{2}(\Omega)$. The nonlinearity $g:[0,3]\times V\times V\to X$ is locally Lipschitz-continuous in a strip along the exact solution for $V=H^{\frac{1}{4}}(\Omega)$.
@@
Figure 6: The convergence rates of ERKC-I methods (in the left panel) and modified ERKC-I methods (in the right panel) for (6.4). The errors are measured at $T=1.4$ in the $L^{\infty}(\Omega)$ norm.